# A Comprehensive Survey of Synthetic Tabular Data Generation

### v2

@@ III-A1 Machine Learning-based Methods
Before the specially designed generative models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models, machine learning (ML) techniques such as Copulas [28], Gaussian Mixture Models (GMMs) [29], and Bayesian networks [83] were widely employed for data generation. These methods rely on explicit parameterization and estimation of joint probability distributions. Additionally, specialized resampling methods were developed, including shallow interpolation-based techniques such as the Synthetic Minority Over-sampling Technique (SMOTE) [30] and its variants (e.g., Borderline-SMOTE [31], SMOTETomek [32], and SMOTE-ENC [33]), as well as Adaptive Synthetic Sampling (ADASYN) [34] and probabilistic methods like Synthpop [35] for data synthesis and interpolation. However, these models struggle with capturing complex data distributions, particularly in high-dimensional settings where accurately learning feature dependencies becomes challenging. Meanwhile, their reliance on predefined assumptions and explicit probability estimation often limits their effectiveness in modeling intricate relationships present in real-world tabular data.
@@ III-A2 VAE-based Method
With the rise of deep learning, generative models specifically designed for tabular data synthesis have gained prominence. Compared to ML learning techniques, deep learning-based generative models, particularly those utilizing VAEs and GANs, have demonstrated significant improvements in generating high-quality synthetic tabular data. Among these, VAE-based methods are shown to achieve superior performance. VAEs learn latent representations by maximizing the Evidence Lower Bound (ELBO) [84], which is defined as:
@@
TVAE [38] adapts VAEs to tabular data by preprocessing and modifying the loss function. VAEM [37] is trained in a two-stage manner such that the first stage provides a more uniform representation of the data to the second stage, thereby sidestepping the problems caused by heterogeneous data. GOGGLE [36] proposes to learn a graph adjacency matrix to model the dependency relationships between different features. The encoder and decoder of the VAE model are both implemented as Graph Neural Networks (GNNs) [85], and the graph adjacent matrix is jointly optimized with the GNNs parameters. Additionally, GAN-based methods have been widely explored in the field of synthetic tabular generation due to their powerful generation capabilities.
@@ III-A3 GAN-based Generators
$\min_{G}\max_{D}\;\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z)))],$ (3)
@@
where the generator $G$ aims to produce a synthetic sample $x_{s}$ that can fool the discriminator $D$, while the discriminator seeks to distinguish real data $x_{e}$ from generated samples. GAN-based methods for tabular data generation are typically categorized into three groups based on their application scenarios: Traditional GAN, Conditional GAN, and Differentially Private GAN.
@@
Traditional GAN. Several studies have extended GANs for tabular data by modifying the model architecture to accommodate categorical variables. TGAN [39] utilizes Long Short-Term Memory (LSTM) [87] with attention mechanisms [88] to generate both continuous and discrete variables feature by feature. MedGAN [40] develops an autoencoder-based generative model for generating high-dimensional electronic health records (EHR). CasTGAN [41] is a generative network framework characterized by multiple generators connected sequentially, where each generator is responsible for generating a single feature. A single discriminator validates the output of all the generators, with training occurring on the final generator in the cascade. CrGAN-Cnet [42] employs GANs for generating airline passenger name records, incorporating Cramer distance [89] and a Cross-Net architecture to improve the model’s handling of missing values. TableGAN [43] introduces an information loss function and a classifier within a GAN framework, utilizing Convolutional Neural Networks (CNNs) [90] for the generator, discriminator, and classifier. IT-GAN [44] proposes a generalized GAN framework for synthetic tabular data generation, combining adversarial training with negative logarithmic density regularization in reversible neural networks, balancing generation quality with privacy leakage. VT-GAN [45] introduces a training-with-shuffling technique to prevent any party from reconstructing the training data from the GAN’s conditional vector, enabling distributed tabular data synthesis in a privacy-preserving manner. FCT-GAN [46] incorporates feature tokenization and Fourier networks to construct a transformer-style generator and discriminator, capturing both local and global dependencies across columns. TabFairGAN [47] uses a two-phase training process: the first phase generates synthetic data similar to the reference dataset, and the second phase modifies the value function to add fairness constraints, ensuring that the synthetic data is both accurate and fair. While these methods enhance synthetic tabular data generation, they lack the ability to conditionally generate data based on specific feature constraints, such as generating health records for female patients.
@@
Conditional GAN. To address the limitations of traditional GAN in controlling the characteristics of synthetic data $\mathcal{D}_{\text{s}}$, Conditional GAN (cGAN) [91] have been widely adopted. The cGAN is an extension of the GAN framework. An additional conditional space $C$ is introduced, which represents the external information coming from the training data. The cGAN framework modifies the model $G$ and $D$, to include the additional conditional space $C$, as follows:
$\begin{split}\min_{G}\max_{D}V(D,G)=&\;\mathbb{E}_{x,c\sim p_{\text{data}}}[\log D(x,c)]\\ &+\mathbb{E}_{z\sim p_{z}(z),c\sim p(c)}[\log(1-D(G(z,c),c))],\end{split}$ (4)
where $(x,c)\in X\times C$ values are sampled from the data distribution $p_{\text{data}}(x,c)$ and $c\in C$ values are sampled from conditional data vectors found in the training data and represented by the density function $p_{y}(c)$.
@@
cGANs [48] incorporate a conditional vector, allowing the generation of data for specific classes, which is particularly useful when dealing with highly imbalanced datasets. This capability is crucial for applications such as initializing datasets in online learning scenarios [92]. CTAB-GAN [49] designs a novel conditional vector that efficiently encodes mixed data types and addresses skewed distributions, mitigating data imbalance and long-tail challenges. OCT-GAN [50] improves utility by designing the generator and discriminator based on Neural Ordinary Differential Equations (NODEs) [93], demonstrating that NODEs have theoretically advantageous properties for generating tabular data. OCT-GAN’s generator includes an ODE layer at the beginning of its architecture to transform the initial input vector into a latent space suitable for the generation process. cWGAN [51] integrates Wasserstein distance [94] into the cGAN framework, leveraging conditional vectors to oversample minority classes for more balanced data generation. CTGAN [38] enhances cGANs by incorporating the PacGAN [95] structure in the discriminator, employing a combination of generator loss, Wasserstein loss with gradient penalty, and a training-by-sampling strategy to handle imbalanced categorical variables. RC-TGAN [52] models relational information between tables by incorporating conditional data from parent rows into the design of child tables. It extends this to model the influence of grandparent rows on grandchild rows, ensuring that these relationships are preserved even when the parent rows fail to transfer relational information.
@@
Differentially Private GAN. To ensure differential privacy during training, the differentially private stochastic gradient descent (DP-SGD) [18] algorithm is widely employed. The core idea is to add calibrated noise to the averaged and clipped per-sample gradients in each iteration. The modified gradient used for updating the model parameters at iteration $t$ is defined as follows:
@@
where $g_{t}^{i}$ denotes the gradient computed on the $i$-th training example in a minibatch of size $L$. Each per-sample gradient is individually clipped using an $\ell_{2}$-norm threshold $C$ to ensure bounded sensitivity:
@@
After clipping, the gradients are averaged and additive Gaussian noise $\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})$ is injected to achieve differential privacy. The parameter $\sigma$ controls the noise scale, balancing the trade-off between privacy and model utility. The identity matrix $\mathbf{I}$ ensures isotropic noise across dimensions. This mechanism ensures that the influence of any single training instance on the model update is limited, thereby enabling formal differential privacy guarantees under standard accounting methods.
DPGAN [18] achieves differential privacy by adding carefully designed noise to the gradients during the learning process. PATE-GAN [19] utilizes the Private Aggregation of Teacher Ensembles (PATE) [96] framework, perturbing the output of multiple teacher discriminators using Laplacian noise [97] before training a student discriminator. However, a key limitation of PATE-GAN [19] is that the student discriminator only learns from synthetic data, which may lead to unreliable feedback. Alternative methods, RDP-CGAN [53] incorporate DP-SGD [18] with Wasserstein loss to improve privacy protection while maintaining model performance. RDP-CGAN[53] employs Rényi Differential Privacy (RDP) [98] accounting, which provides tighter bounds on privacy costs and enhances the privacy-utility trade-off. CTAB-GAN+ [54] builds upon these methods by incorporating RDP-based privacy accounting, similar to DP-WGAN [99], but using a single discriminator instead of multiple ones, as seen in PATE-GAN [19] and GS-WGAN [100]. By leveraging Was+GP loss, CTAB-GAN+ [54] intrinsically constrains the gradient norm, eliminating the need for weight clipping and resulting in more stable training for differentially private GANs.
@@ III-B Diffusion Model Methods
Model Year Latent Space Venue Domain
DDPM-based TabDDPM [14] 2023 No ICML Generic
CoDi [55] 2023 No ICML Generic
AutoDiff [57] 2023 Yes SyntheticData4ML Generic
RelDDPM [58] 2024 No Proc. ACM Manag. Data Generic
CTSyn [56] 2025 Yes ICLR Generic
TABDIFF [15] 2025 No ICLR Generic
TabDDPM-HER [59] 2023 No arXiv Healthcare
DPM-HER [60] 2023 No SyntheticData4ML Healthcare
FLEXGEN-EHR [61] 2024 Yes ICLR Healthcare
EntTabDiff [62] 2023 Yes ICAIF Finance
FinDiff [63] 2024 No ICAIF Finance
Imb-FinDiff [64] 2024 Yes ICAIF Finance
Score-based SOS [65] 2022 No KDD Generic
MissDiff [66] 2022 No TRL @ NeurIPS Generic
CSDI_T [67] 2023 No SPIGM @ ICML Generic
STaSY [68] 2023 No ICLR Generic
Forest-VP/Forest-Flow [69] 2024 No PMLR Generic
TABSYN [70] 2024 Yes ICLR Generic
Although traditional deep learning methods such as VAEs and GANs have made notable progress in synthetic tabular data generation, these models often encounter challenges, including training instability, mode collapse, and poor representation of multimodal distributions. These limitations hinder the generation of high-quality synthetic tabular data. In contrast, diffusion models effectively address these challenges through their specialized denoising mechanism, which has garnered significant research interest. In the domain of tabular data, diffusion models have begun to exhibit advantages similar to those of GANs and VAEs, achieving notable performance improvements. These models demonstrate strong potential in overcoming the unique challenges associated with tabular data modeling, highlighting their effectiveness in generating high-quality, diverse synthetic data. The current mainstream tabular diffusion generation models include DDPM-based Method, and Score-based Method as shown in Table II.
@@ III-B1 DDPM-based Method
Diffusion models [13, 101] are likelihood-based generative models that synthesize data by learning to reverse a gradual noising process. They model data generation as a Markovian process consisting of two stages: a forward (diffusion) process, which progressively adds noise to data, and a reverse (denoising) process, which reconstructs data from noise. Given an initial data point $x_{0}\sim q(x_{0})$, the forward process defines a series of latent variables $x_{1},x_{2},\ldots,x_{T}$ by iteratively adding noise sampling from the predefined distributions $q(x_{t}|x_{t-1})$ with variances $\{\beta_{1},...,\beta_{T}\}$:
@@
$\begin{split}\log q(x_{0})\geq\mathbb{E}_{q(x_{0:T})}\Big{[}&\log p_{\theta}(x_{0}\mid x_{1})-\mathrm{KL}(q(x_{T}\mid x_{0})\,\|\,q(x_{T}))\\ &-\sum_{t=2}^{T}\mathrm{KL}(q(x_{t-1}\mid x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}\mid x_{t}))\Big{]}.\end{split}$ (9)
@@
Due to the inherent heterogeneity of tabular data, existing Denoising Diffusion Probabilistic Models (DDPMs) face challenges in direct application to tabular data. To address this, researchers have proposed modifications to DDPMs to better accommodate the structural characteristics of tabular data. TabDDPM [14] extends DDPM to tabular data and demonstrates its superiority over existing GAN/VAE alternatives, which is consistent with the advantages of diffusion models in other fields. TABDIFF [15] develops a joint continuous-time diffusion process for numerical and categorical data and proposes feature-wise learnable diffusion processes to counter the diversity of different feature distributions. CoDi [55] uses two diffusion models to handle continuous and discrete variables, respectively. During the training process, these two diffusion models mutually read conditions and constrain each other, solving the problem of modeling discrete features. CTSyn [56] develops a unified aggregator that tokenizes and embeds heterogeneous table rows, projects them into a unified latent space, and facilitates the training of models across tabular data. AutoDiff [57] leverages autoencoders to encode tabular data into a latent space, where the diffusion process is applied. This approach effectively addresses challenges associated with heterogeneous features, mixed data types, and feature correlations, enabling more accurate and structured synthetic tabular data generation. RelDDPM [58] leverages diffusion models to first learn an unconditional generative model. Subsequently, RelDDPM [58] introduced lightweight controllers to guide the unconditional generative model in generating synthetic data that satisfies different conditions.
@@
In addition to employing DDPM-based methods for general-purpose synthetic tabular data generation, several studies have explored their application in specific domains—most notably, healthcare. For instance, TabDDPM-EHR [59] applies TabDDPM to the generation of electronic health records (EHR), demonstrating its ability to synthesize high-quality EHR data while preserving the statistical properties of the original dataset. DPM-EHR [60] leverages DDPM to generate EHR data consisting of mixed-type variables, including numerical, binary, and categorical features. Furthermore, FLEXGEN-EHR [61] introduces an optimal transformation module that aligns heterogeneous EHRs into a common feature space, while effectively addressing missing modalities within a unified learning framework. In the finance domain, diffusion-based models have also shown promising results. EntTabDiff [62] generates synthetic entities by learning the distribution of entity-level features and subsequently synthesizes tabular data conditioned on these entities. FinDiff [63] is designed to produce high-fidelity synthetic financial tabular data suitable for various downstream regulatory tasks. Building upon this, Imb-FinDiff [64] further addresses the challenge of class imbalance by focusing on the generation of samples from minority categories within long-tail distributions, thereby enhancing the representativeness and utility of synthetic financial datasets.

### v3

@@ III-A1 Machine Learning-based Methods
Before the specially designed generative models such as Variational Autoencoders (VAEs), and Generative Adversarial Networks (GANs), machine learning (ML) techniques such as Copulas [33], Gaussian Mixture Models (GMMs) [34], and Bayesian networks [88] were widely employed for data generation. These methods rely on explicit parameterization and estimation of joint probability distributions. Additionally, specialized resampling methods were developed, including shallow interpolation-based techniques such as the Synthetic Minority Over-sampling Technique (SMOTE) [35] and its variants (e.g., Borderline-SMOTE [36], SMOTETomek [37], and SMOTE-ENC [38]), as well as Adaptive Synthetic Sampling (ADASYN) [39] and probabilistic methods like Synthpop [40] for data synthesis and interpolation. However, these models struggle with capturing complex data distributions, particularly in high-dimensional settings where accurately learning feature dependencies becomes challenging. Meanwhile, their reliance on predefined assumptions and explicit probability estimation often limits their effectiveness in modeling intricate relationships present in real-world tabular data.
@@ III-A2 VAE-based Method
With the rise of deep learning, generative models specifically designed for tabular data synthesis have gained prominence. Compared to ML learning techniques, deep learning-based generative models, such as VAE, further improve the quality of synthetic tabular data by learning latent representations by maximizing the Evidence Lower Bound (ELBO) [89], which is defined as:
@@
TVAE [43] adapts VAEs to tabular data by transforming categorical features into numerical ones using one-hot encoding. VAEM [42] is trained in a two-stage manner, the first stage provides a more uniform representation of the data to the second stage, thereby avoiding the problems caused by heterogeneous data. GOGGLE [41] introduces Graph Neural Networks (GNN) [90] into the encoder and decoder of VAE, and learns a graph adjacency matrix to capture inter-column dependencies. Although VAE-based Methods have been widely used in the field of synthetic tabular generation, the generated tends towards averaging and loss of details.
@@ III-A3 GAN-based Generators
$\min_{G}\max_{D}\;\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))],$ (3)
@@
where the generator $G$ aims to produce a synthetic sample $x_{s}$ that can fool the discriminator $D$, while the discriminator seeks to distinguish real data $x_{e}$ from generated samples. GAN-based methods for tabular data generation are typically categorized into three groups: Traditional GAN, Conditional GAN, and Differentially Private GAN.
@@
Traditional GAN. Several studies have extended GANs for tabular data by modifying the model architecture to accommodate categorical features. TGAN [44] utilizes Long Short-Term Memory (LSTM) [92] with attention mechanisms [93] to generate both numerical and categorical features one by one. MedGAN [45] develops an autoencoder-based generative model for generating high-dimensional electronic health records (EHR). CasTGAN [46] is a generative network framework characterized by multiple generators connected sequentially, where each generator is responsible for generating a feature, and one discriminator validates the output of all the generators. CrGAN-Cnet [47] employs GANs for generating airline passenger name records, incorporating Cramer distance [94] and a Cross-Net architecture to improve the model’s handling of missing values. TableGAN [48] uses Convolutional Neural Networks (CNNs) [95] for both the generator and discriminator, when the tabular data includes a label column, a prediction loss is added to the generator to better improve the correlation between the label column and the other columns. IT-GAN [49] proposes a generalized GAN framework for synthetic tabular data generation, combining adversarial training with negative logarithmic density regularization in reversible neural networks, and balances the generation quality with privacy protection leakage. VT-GAN [50] introduces a training-with-shuffling technique to prevent any party from reconstructing the training data from the GAN’s conditional vector, enabling distributed tabular data synthesis in a privacy protection manner. FCT-GAN [51] incorporates feature tokenization and Fourier networks to construct a transformer-style generator and discriminator, capturing both local and global inter-column dependencies. TabFairGAN [52] uses a two-phase training process: the first phase generates synthetic data similar to the reference dataset, and the second phase modifies the value function to add fairness constraints, ensuring that the synthetic data is both accurate and fair. While these methods enhance synthetic tabular data generation, they lack the ability to conditionally generate data based on specific feature constraints, such as generating health records for female patients.
@@
Conditional GAN. To address the limitations of traditional GAN in controlling the features of synthetic data $\mathcal{D}_{\text{s}}$, Conditional GAN [96] have been widely adopted,which is an extension of the GAN framework. An additional conditional space $C$ is introduced, which represents the external information coming from the training data. The Conditional GAN framework modifies the model $G$ and $D$, to include the additional conditional space $C$, as follows:
$\begin{split}\min_{G}\max_{D}V(D,G)&=\;\mathbb{E}_{x,c\sim p_{\text{data}}(x,c)}[\log D(x,c)]\\ &+\mathbb{E}_{z\sim p_{z}(z),c\sim p_{y}(c)}[\log(1-D(G(z,c),c))],\end{split}$ (4)
where $(x,c)\in X\times C$ are sampled from the data distribution $p_{\text{data}}(x,c)$.
@@
cGANs [53] incorporates a conditional vector, allowing the generation of data for specific classes, which is particularly useful when dealing with highly imbalanced datasets. This capability is crucial for applications such as initializing datasets in online learning scenarios [97]. CTAB-GAN [54] designs a novel conditional vector that efficiently encodes mixed data types and addresses skewed distributions, mitigating data imbalance and long-tail challenges. OCT-GAN [55] designs the generator and discriminator based on Neural Ordinary Differential Equations (NODEs) [98], demonstrating that NODEs have theoretically advantageous properties for synthetic tabular data generation. OCT-GAN’s generator includes an ODE layer at the beginning of its architecture to transform the initial input vector into a latent space suitable for the generation process. cWGAN [56] integrates Wasserstein distance [99] into the Conditional GAN framework, leveraging conditional vectors to oversample minority classes for more balanced data generation. CTGAN [43] enhances cGANs [53] by incorporating the PacGAN [100] structure in the discriminator, employing a combination of generator loss, Wasserstein loss with gradient penalty, and a training-by-sampling strategy to handle imbalanced categorical features. RC-TGAN [57] is proposed to handle complex multi-tabular datasets by conditioning the generation of the child tabular data on the previously generated parent rows, specifically using the features of the parent. This method transfers relationship information from the parent tabular data to the child tabular data.
@@
Differentially Private GAN. To ensure differential privacy protection during training, the differentially private stochastic gradient descent (DP-SGD) [18] algorithm is widely employed. The core idea is to add calibrated noise to the averaged and clipped per-sample gradients in each iteration. The modified gradient used for updating the model parameters at iteration $t$ is defined as follows:
@@
where $g_{t}^{i}$ denotes the gradient computed on the $i$-th training sample in a minibatch of size $L$. Each per-sample gradient is individually clipped using an $\ell_{2}$-norm threshold $C$ to ensure bounded sensitivity:
@@
After clipping, the gradients are averaged and additive Gaussian noise $\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})$ is injected to achieve differential privacy protection. The parameter $\sigma$ controls the noise scale, balancing the trade-off between privacy protection and model utility. The identity matrix $\mathbf{I}$ ensures isotropic noise across dimensions. This mechanism ensures that the influence of any single training sample on the model update is limited, thereby making it impossible for attackers to infer specific individual information even if they obtain model output.
DPGAN [18] achieves differential privacy protection by adding carefully designed noise to the gradients during the learning process. PATE-GAN [19] utilizes the Private Aggregation of Teacher Ensembles (PATE) [101] framework, perturbing the output of multiple teacher discriminators using Laplacian noise [102] before training a student discriminator. However, a key limitation of PATE-GAN [19] is that the student discriminator only learns from synthetic data, which may lead to unreliable feedback. RDP-CGAN [58] builds upon convolutional autoencoders and generative adversarial networks, incorporating Rényi Differential Privacy (RDP) [103] to provide tighter bounds on privacy costs and improve the privacy-utility trade-off. CTAB-GAN+ [59] builds upon existing methods by incorporating RDP-based privacy accounting, similar to DP-WGAN [104]. Furthermore, by leveraging the Was+GP loss, CTAB-GAN+ effectively constrains the gradient norm, eliminating the need for weight clipping and resulting in more stable training for differentially private GANs.
@@ III-B Diffusion Model Methods
Model Year Latent Space Venue
DDPM-based TabDDPM [14] 2023 No ICML
CoDi [60] 2023 No ICML
AutoDiff [62] 2023 Yes SyntheticData4ML
RelDDPM [63] 2024 No Proc. ACM Manag. Data
CTSyn [61] 2025 Yes ICLR
TABDIFF [15] 2025 No ICLR
TabDDPM-HER [64] 2023 No arXiv
DPM-HER [66] 2023 No SyntheticData4ML
FLEXGEN-EHR [68] 2024 Yes ICLR
EntTabDiff [69] 2023 Yes ICAIF
FinDiff [65] 2024 No ICAIF
Imb-FinDiff [67] 2024 Yes ICAIF
Score-based SOS [70] 2022 No KDD
MissDiff [71] 2022 No TRL @ NeurIPS
CSDI_T [72] 2023 No SPIGM @ ICML
STaSY [73] 2023 No ICLR
Forest-VP/Forest-Flow [74] 2024 No PMLR
TABSYN [75] 2024 Yes ICLR
Figure 4: The Training and sample of Diffusion Model Methods.
Although GANs have made notable progress in synthetic tabular data generation, these models often encounter challenges, including training instability, mode collapse, and poor representation of multimodal distributions, which hinder the generation of high-quality synthetic tabular data. In contrast, diffusion models effectively address these challenges through their specialized denoising mechanism, which gains significant research interest. The training and sampling process of diffusion model methods is shown in Figure 4. In the domain of synthetic tabular data generation, diffusion models have begun to exhibit advantages similar to those of GANs and VAEs, achieving notable performance improvements. These models demonstrate strong potential in overcoming the unique challenges associated with complex tabular data distributions modeling, highlighting their effectiveness in generating high-quality, diverse synthetic data. The current mainstream tabular diffusion generation models include DDPM-based Method, and Score-based Method as shown in Table II.
@@ III-B1 DDPM-based Method
Diffusion models [12, 105] are likelihood-based generative models that synthesize data by learning to reverse a gradual noising process, which model data generation as a Markovian process consisting of two stages: a forward (diffusion) process, which progressively adds noise to data, and a reverse (denoising) process, which reconstructs data from noise. Given an initial data point $x_{0}\sim q(x_{0})$, the forward process defines a series of latent features $x_{1},x_{2},\ldots,x_{T}$ by iteratively adding noise sampling from the predefined distributions $q(x_{t}|x_{t-1})$ with variances $\{\beta_{1},...,\beta_{T}\}$:
@@
$\begin{split}\log q(x_{0})\geq&\\ \mathbb{E}_{q(x_{0:T})}\Big{[}&\log p_{\theta}(x_{0}\mid x_{1})-\mathrm{KL}(q(x_{T}\mid x_{0})\,\|\,q(x_{T}))\\ &-\sum_{t=2}^{T}\mathrm{KL}(q(x_{t-1}\mid x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}\mid x_{t}))\Big{]}.\end{split}$ (9)
@@
Due to the inherent heterogeneity of tabular data, existing Denoising Diffusion Probabilistic Models (DDPMs) face challenges in adapting to tabular data. To address this, researchers modify DDPMs to better accommodate the structural characteristics of tabular data. TabDDPM [14] extends DDPM to tabular data and demonstrates its superiority over existing GAN/VAE alternatives, which is consistent with the advantages of diffusion models in other fields. TABDIFF [15] develops a joint continuous-time diffusion process for numerical and categorical data and proposes feature-wise learnable diffusion processes to counter the diversity of different feature distributions. CoDi [60] uses two diffusion models to handle numerical and categorical features, respectively. During the training process, these two diffusion models mutually read conditions and constrain each other, solving the problem of modeling categorical features. CTSyn [61] develops a unified aggregator that tokenizes and embeds heterogeneous tabular rows to a unified latent space, and facilitates the training of models across tabular data. AutoDiff [62] leverages autoencoders to encode tabular data into a latent space, where the diffusion process is applied. This approach effectively addresses challenges associated with heterogeneous features, mixed data types, and complex inter-column dependencies, enabling more accurate and structured synthetic tabular data generation. RelDDPM [63] leverages diffusion models to first learn an unconditional generative model. Subsequently, RelDDPM [63] introduced lightweight controllers to guide the unconditional generative model in generating synthetic data that satisfies different conditions.
@@
In addition to employing DDPM-based methods for general-purpose synthetic tabular data generation, several studies have explored their application in specific domains—most notably, healthcare. For instance, TabDDPM-EHR [64] applies TabDDPM to the generation of electronic health records (EHR), demonstrating its ability to synthesize high-quality EHR data while preserving the statistical properties of the original dataset. DPM-EHR [66] leverages DDPM to generate EHR data consisting of mixed-type features, including numerical, binary, and categorical features. Furthermore, FLEXGEN-EHR [68] introduces an optimal transformation module that aligns heterogeneous EHRs into a common feature space, while effectively addressing the promplem of missing values within a unified learning framework. In the finance domain, diffusion-based models have also shown promising results. EntTabDiff [69] generates synthetic entities by learning the distribution of entity-level features and subsequently synthesizes tabular data conditioned on these entities. FinDiff [65] is designed to produce high-fidelity synthetic financial tabular data suitable for various downstream tasks. Building upon this, Imb-FinDiff [67] further addresses the challenge of class imbalance by focusing on the generation of samples from minority categories within long-tail distributions, thereby enhancing the representativeness and utility of synthetic financial datasets.